OpenStudy (anonymous):

5x - 7 = 3(x - 2) + 3 please show me the steps so i can get through the rest of my problems nd explain

OpenStudy (anonymous):

do the distrubutive property on 3(x-2)= 3x-6+3 and ull have 5x-7=3x-6+3. What you do to one side do to the other.. if your solving for x u do -6+3=-3. u add 7 to both sides and and have 5x=3x-3. subtract 3x to 5x then hav 2x=-3 then divid by two to get -o.6 repeted

OpenStudy (anonymous):

abbie2016 is wrong. 5x-7=3x-6+3. Isolate the variable. Subtract 3x and add 7 to both sides. You get 2x=4. So x=2.

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

ndog33

OpenStudy (anonymous):

5x - 7 = 3(x - 2) + 3 5x - 7 = 3x - 6 + 3................ any brackets taken out by multiplication 5x -3x= -6+3 +7 ...........bring x's to one side , numbers to the other side, sign changes if they cross sides. 2x = 4 x=2.

OpenStudy (anonymous):

lol did you good answer abbie

OpenStudy (anonymous):

i wasn't tryin too